# Conditioning diffusion models by explicit forward-backward bridging

### v1

@@ 5 Related works
In the past few years, several methods have been proposed to perform conditional sampling within diffusion models. The first way to do so was proposed alongside score-based denoising diffusion models in the landmark work by Song et al. (2021). They rely on learning an approximate conditioning drift model after the fact, adding a layer of approximate learning on top of the unconditional one. This type of idea can also be found in Shi et al. (2022) who train a Schrödinger bridge on the joint distribution of the latent variable $X$ and the observed $Y$ when the model $\pi(y\mid x)$ is available. Closer to our work, Wu et al. (2023) propose using a twisted particle filter (Whiteley and Lee, 2014) to efficiently bridge between the reference distribution and the conditional $\pi(x\mid y)$ at the cost of additional expensive computations in the procedure. Finally, Trippe et al. (2023), as well as the contemporaneous Dou and Song (2024), express the problem of conditioning as a bridge for simulated observations too. In fact, Trippe et al. (2023) is exactly given as a special case of Section 3.2 where the new path is proposed independently from the previous one (i.e., in our formulation, $q\equiv\mathbb{F}(\cdot\mid y_{0})$), and the resulting sample is accepted unconditionally (i.e., their method is equivalent to setting $\alpha=1$), ignoring the fact that two $Y$ paths sampled under the forward dynamics (the prior) may have different marginal likelihoods under the backward denoising model. This means that their method is inherently biased and does not asymptotically target the right distribution, despite their consistency result, even under the assumption $\pi_{T}\equiv\pi_{\mathrm{ref}}$. The resulting bias is reflected in our experiments. This defect is shared by Dou and Song (2024). Additional background details can be found in Appendix A.
@@ 6 Conclusion
We have presented a novel MCMC-based method for sampling from conditional distributions expressed as diffusion models and Schrödinger bridges. Our method relies on formulating the sampling procedure as an inference problem on missing observation data, thanks to which we can apply classical particle MCMC algorithms (Andrieu et al., 2010) to perform (asymptotically) exact sampling within the model. We have demonstrated the performance and correctness of our method on simulated and real datasets, for inpainting and super-resolution problems alike.
@@
While our method improves on the pre-existing methodology, in particular Trippe et al. (2023) on which we generalise, several questions remain open. In particular, (i) while we did not directly report on this, we observed that the distribution of the particles in the backward filtering pass contracts at the terminal time (Trippe et al., 2023; Wu et al., 2023, this is true for all methods considered here, including), and may consequently increase the autocorrelation of our MCMC chains, explaining their lack of diversity despite the overall improvement in quality. Because this appears throughout, it seems to be a feature of the model decomposition rather than a feature of Gibbs-CSMC or PMCMC. (ii) We also notice that the variance of the log-likelihood estimation in PMCMC can be large, and hence, its acceptance rate can be unstable, also making the calibration of its step-size difficult. While this is detrimental to our approach, it also explains the poor posterior coverage of Trippe et al. (2023) who completely ignore this step.
@@
Consequently, to obtain independent samples with adequate posterior coverage, a pragmatic compromise is likely to repeatedly use Trippe et al. (2023) as initialisation and then correct the posterior sample by running our methods for a few iterations. Nevertheless, solving both these issues is of great interest, and is likely feasible within the framework we outlined in this article by a more careful consideration of the structure of the augmented model. We leave these questions for future work.
Individual Contributions
The original idea and methodology are due to AC and subsequently refined jointly by AC and ZZ. Implementation and experimental evaluation are mostly due to ZZ with help and inputs from AC. Writing was primarily done by AC with substantial help from ZZ and inputs from TS. All authors edited and validated the final manuscript.
Acknowledgement
@@
This work was partially supported by the Kjell och Märta Beijer Foundation, and Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. The computations were enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre. Adrien Corenflos and Simo Särkkä were partially supported by the Research Council of Finland. Adrien Corenflos also acknowledges the financial support provided by UKRI for OCEAN (a 2023-2029 ERC Synergy grant co-sponsored by UKRI).
@@ A.1.1 Denoising diffusion models
Denoising diffusion models (Song et al., 2021) have recently received a lot of attention as a general-purpose sampler for generative modelling (Song et al., 2021) or statistical inference (Vargas et al., 2023). At the core, given a target distribution $\pi_{0}$ and a time-homogeneous stochastic differential equation (SDE)
@@
with stationary distribution $\pi_{\mathrm{ref}}$, they sample from $\pi_{0}$ by “denoising” (8) by means of Doob’s $h$-transform (Rogers and Williams, 2000). Formally, let $\pi_{t}$ be the distribution of $X_{t}$ under (8), and assume that we know how to sample from $\pi_{T}$ for some $T>0$. We can obtain approximate samples from $\pi_{0}$ by sampling from
@@
When $T\gg 1$, and under ergodicity guarantees (Meyn and Tweedie, 2009), $\pi_{T}\approx\pi_{\mathrm{ref}}$, so that if (8) is chosen such that $\pi_{\mathrm{ref}}$ is easy to sample from, then the only remaining blocker is computing the score $\nabla\log\pi_{T-t}$. However, obtaining an expression for $\pi_{t}$ (and therefore for $\nabla\log\pi_{T-t}$) is in general not possible, and denoising diffusion models usually rely on score matching (Hyvärinen, 2005), whereby an approximation $s_{\theta^{*}}(t,x_{t})$ to $\nabla\log\pi_{t}$ is learnt under samples by minimising the loss function
@@
with $0=t_{0}<t_{1}<\cdots<t_{K}=T$ being a given integration grid used also at evaluation time to simulate from (9). When (8) is chosen, as is often the case, to be linear (for example, as an Ornstein–Uhlenbeck process, Uhlenbeck and Ornstein, 1930), $\pi_{t}(\cdot\mid X_{0})$ can be computed and sampled from exactly so that the first version of (11) incurs no error other than the sampling noise. When it is not, as in the next section, the second version is more amenable to computations (and is also more numerically stable), at the cost of storing trajectories under $\mathbb{P}$.
@@ A.1.2 Schrödinger bridges
Schrödinger bridges (Léonard, 2014; Chen et al., 2021) constitute a class of bridges which aim at minimising the total energy expense to move from $\pi_{0}$ to $\pi_{T}$ and vice versa. Given a reference path measure $\mathbb{Q}$ on the set of continuous paths $\mathcal{C}([0,T],\mathbb{R}^{d})\eqqcolon\mathcal{C}$, the Schrödinger bridge solution $\mathbb{P}^{\mathrm{SB}}$ is defined as the (unique) solution to
@@
where $\mathcal{P}(\mathcal{C})$ denotes all the path measures on $\mathcal{C}$. In general, solving (12) directly is not possible, and instead, one needs to rely on alternative ways of solving the Schrödinger bridge. The two main methods to do so are given by iterative proportional fitting (IPF, De Bortoli et al., 2021) and Schrödinger bridge matching (SBM, Shi et al., 2023) which we review below.
@@ A.2.1 Conditional score matching
Score-based generative models can readily generate conditional samples from pre-trained unconditional models by plugging $\nabla_{x}\log\pi(y\mid x)$ into the drift of (9). However, it is in general hard to know $\pi(y\mid x)$ or its score, for example, in image restoration/classification applications (Song et al., 2021). Due to this, one often needs to approximate a separate time-varying model for $\pi_{t}(y\mid x_{t})$ or make domain-based heuristics, but this inevitably incurs additional errors for the conditional sampling.
@@ A.2.2 Conditional Schrödinger bridges
Song et al. (2021) which requires learning or approximating the likelihood $\pi_{t}(y\mid x_{t})$. By contrast, Shi et al. (2022) propose a method to learn a conditional sampler only requiring being able to sample artificial observations $\pi(y\mid x)$. The idea is built upon constructing a Schrödinger bridge (see Section A.1.2) for the augmented distribution $\pi(x,y)$, namely
@@
with initial distribution $X_{0},Y_{0}\sim\pi(x,y)$ and an arbitrary drift $a$. In Equation (14), $\pi_{\mathrm{ref}}(y\mid x)$ refers to any informative approximation to $\pi(x\mid y)$ that is easy to sample from. We can then apply the methodology introduced in Section A.1.2 to solve Equation (14).
@@ A.3 Twisted diffusion samplers
for which the conditional distribution of $U_{T}\mid y$ is $\pi(\cdot\mid y)$. This corresponds to a state-space model with a final observation $Y=y$, which can therefore be sampled by running the reverse diffusion (16) and then correcting the samples via importance sampling at the final time step. This is however well-known to be inefficient (Whiteley and Lee, 2014) as the unconditional samples $U_{T}$ are unlikely to be distributed according to $\pi(\cdot\mid y_{T})$.
In order to efficiently sample from $\pi(\cdot\mid y)$, Wu et al. (2023) proposed to form the twisted model
@@
which marginally recovers $\pi(U_{T}\mid y)$. Once an approximation $\tilde{p}(u_{t_{k+1}}\mid y,u_{t_{k}})$ to the true conditional process $p(u_{t_{k+1}}\mid y,u_{t_{k}})$ is chosen, the twisted model (17) can then be sampled from using particle filtering, with a proposal $\tilde{p}(u_{t_{k+1}}\mid y,u_{t_{k}})$ and the corresponding importance weights as given in (18), see also Wu et al. (2023); Whiteley and Lee (2014) for more details. While principled, this approach presents at least the following two issues: (i) it requires the likelihood model $\pi(y\mid x)$ to be known, which is not the case in most applications of interest, (ii) and it is computationally intensive as forming the approximation $\tilde{p}$ requires the denoising score function to be differentiated (Wu et al., 2023, Section 3.2).
@@ A.4 Pathwise conditioning
The idea of conditioning on the path of a diffusion process has been explored in the context of diffusion models by Trippe et al. (2023) for scaffolding and by Dou and Song (2024, concurrent with our paper) for the slighly more general setting of inverse problems. The idea is that, given an observation $Y=y$, and provided that the forward noising diffusion for $X_{t},Y_{t}$ is separable, one can form a sequence of observations $Y_{0:K}$ and then sample from the conditional distribution $\pi(X_{0:K}\mid Y_{0:K})$ by running the reverse diffusion (5) conditioned on $Y_{0:K}$ for instance by using a particle filter as in Wu et al. (2023). Beside the need for a separable noising SDE, the method is asymptotically consistent under three assumptions:
@@
Appendix B Experiment settings

### v2

@@ 5 RELATED WORKS
In the past few years, several methods have been proposed to perform conditional sampling within diffusion models (see e.g., a review in Zhao et al.,, 2024). The first way to do so was proposed alongside score-based denoising diffusion models in the landmark work by Song et al., (2021). They rely on learning an approximate conditioning drift model after the fact, adding a layer of approximate learning on top of the unconditional one (see typical approximations in Chung et al.,, 2023; Song et al.,, 2023; Mardani et al.,, 2024; Denker et al.,, 2024). This type of idea can also be found in Shi et al., (2022) who train a Schrödinger bridge on the joint distribution of the latent variable $X$ and the observed $Y$ when the model $\pi(y\mid x)$ is available. Closer to our work, Wu et al., (2023) propose using a twisted particle filter (Whiteley and Lee,, 2014) to efficiently bridge between the reference distribution and the conditional $\pi(x\mid y)$ at the cost of additional expensive computations in the procedure. Finally, Trippe et al., (2023), as well as the contemporaneous Dou and Song, (2024), express the problem of conditioning as a bridge for simulated observations too. In fact, Trippe et al., (2023) is exactly given as a special case of Section 3.2 where the new path is proposed independently from the previous one (i.e., in our formulation, $q\equiv\mathbb{F}(\cdot\mid y_{0})$), and the resulting sample is accepted unconditionally (i.e., their method is equivalent to setting $\alpha=1$), ignoring the fact that two $Y$ paths sampled under the forward dynamics (the prior) may have different marginal likelihoods under the backward denoising model. This means that their method is inherently biased and does not asymptotically target the right distribution, despite their consistency result, even under the assumption $\pi_{T}\equiv\pi_{\mathrm{ref}}$. The resulting bias is reflected in our experiments. This defect is shared by Dou and Song, (2024). Additional background details can be found in Appendix A.
@@ 6 CONCLUSION
We have presented a novel MCMC-based method for sampling from conditional distributions expressed as diffusion models and Schrödinger bridges. Our method relies on formulating the sampling procedure as an inference problem on missing observation data, thanks to which we can apply classical particle MCMC algorithms (Andrieu et al.,, 2010) to perform (asymptotically) exact sampling within the model. We have demonstrated the performance and correctness of our method on simulated and real datasets, for inpainting and super-resolution problems alike.
One iteration of CSMC (or the pseudo-marginal version of the algorithm) has exactly the same computational cost as its unconditional version for the same number of particles. This means that the additional cost is fully encapsulated by the number of iterations required (see, e.g., Figure 8) to achieve convergence. Consequently, to obtain independent samples with adequate posterior coverage, a pragmatic compromise is to repeatedly use Trippe et al., (2023) as initialisation and then correct the posterior sample by running our methods for a few iterations, preferably the CSMC version due to its improved stability.
@@
While our method improves on the pre-existing methodology, in particular Trippe et al., (2023) on which we generalise, several questions remain open. In particular, (i) we observed that the distribution of the particles in the backward filtering pass contracts (see, Appendix F) at the terminal time (Trippe et al.,, 2023; Wu et al.,, 2023, this is true for all methods considered here, including), and may consequently increase the autocorrelation of our MCMC chains, explaining their lack of diversity despite the overall improvement in quality. Because this appears throughout, it seems to be a feature of the model decomposition rather than a feature of Gibbs-CSMC or PMCMC. (ii) We also notice that the variance of the log-likelihood estimation in PMCMC can be large, and hence, its acceptance rate can be unstable, also making the calibration of its step size $\delta$ difficult. While this is detrimental to our approach, it also explains the empirical poor posterior coverage of Trippe et al., (2023) who completely ignore this step. Because of this, we recommend using the Gibbs-CSMC approach above PMCMC, despite the latter’s occasional better experimental performance.
@@
Nevertheless, solving both these two issues of the framework is of great interest, and is likely feasible within the framework we outlined in this article by a more careful consideration of the structure of the augmented model. We leave these questions for future work.
Finally, we remark that our method can straightforwardly extend to a wider class of inverse problems (see, e.g., surveys in Daras et al.,, 2024; Huang et al., 2024b, ; Huang et al., 2024a, ). Notably, our method can, without dedicated training, solve all inverse problems with linear Gaussian likelihood models which can be represented as inpainting problems (Kawar et al.,, 2021). Extending the method to nonlinear likelihood models, which our work does not support, would also make the method of more use to practitioners, and, as a consequence, it is a very important avenue of future work.
Acknowledgements
@@
The original idea and methodology are due to AC and subsequently refined jointly by AC and ZZ. Implementation and experimental evaluation are mostly due to ZZ with help and inputs from AC. Writing was primarily done by AC with substantial help from ZZ and inputs from TS. All authors edited and validated the final manuscript. This work was partially supported by the Kjell och Märta Beijer Foundation, Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, the project Deep probabilistic regression – new models and learning algorithms (contract number: 2021-04301), funded by the Swedish Research Council, and by the Research Council of Finland. The computations were enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre. Adrien Corenflos and Simo Särkkä were partially supported by the Research Council of Finland. Adrien Corenflos also acknowledges the financial support provided by UKRI for OCEAN (a 2023-2029 ERC Synergy grant co-sponsored by UKRI).
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
@@ A.1.1 Denoising diffusion models
Denoising diffusion models (Song et al.,, 2021) have recently received a lot of attention as a general-purpose sampler for generative modelling (Song et al.,, 2021) or statistical inference (Vargas et al.,, 2023). At the core, given a target distribution $\pi_{0}$ and a time-homogeneous stochastic differential equation (SDE)
@@
with stationary distribution $\pi_{\mathrm{ref}}$, they sample from $\pi_{0}$ by “denoising” (8) by means of Doob’s $h$-transform (Rogers and Williams,, 2000). Formally, let $\pi_{t}$ be the distribution of $X_{t}$ under (8), and assume that we know how to sample from $\pi_{T}$ for some $T>0$. We can obtain approximate samples from $\pi_{0}$ by sampling from
@@
When $T\gg 1$, and under ergodicity guarantees (Meyn and Tweedie,, 2009), $\pi_{T}\approx\pi_{\mathrm{ref}}$, so that if (8) is chosen such that $\pi_{\mathrm{ref}}$ is easy to sample from, then the only remaining blocker is computing the score $\nabla\log\pi_{T-t}$. However, obtaining an expression for $\pi_{t}$ (and therefore for $\nabla\log\pi_{T-t}$) is in general not possible, and denoising diffusion models usually rely on score matching (Hyvärinen,, 2005), whereby an approximation $s_{\theta^{*}}(t,x_{t})$ to $\nabla\log\pi_{t}$ is learnt under samples by minimising the loss function
@@
with $0=t_{0}<t_{1}<\cdots<t_{K}=T$ being a given integration grid used also at evaluation time to simulate from (9). When (8) is chosen, as is often the case, to be linear (for example, as an Ornstein–Uhlenbeck process, Uhlenbeck and Ornstein,, 1930), $\pi_{t}(\cdot\mid X_{0})$ can be computed and sampled from exactly so that the first version of (11) incurs no error other than the sampling noise. When it is not, as in the next section, the second version is more amenable to computations (and is also more numerically stable), at the cost of storing trajectories under $\mathbb{P}$.
@@ A.1.2 Schrödinger bridges
Schrödinger bridges (Léonard,, 2014; Chen et al.,, 2021) constitute a class of bridges which aim at minimising the total energy expense to move from $\pi_{0}$ to $\pi_{T}$ and vice versa. Given a reference path measure $\mathbb{Q}$ on the set of continuous paths $\mathcal{C}([0,T],\mathbb{R}^{d})\eqqcolon\mathcal{C}$, the Schrödinger bridge solution $\mathbb{P}^{\mathrm{SB}}$ is defined as the (unique) solution to
@@
where $\mathcal{P}(\mathcal{C})$ denotes all the path measures on $\mathcal{C}$. In general, solving (12) directly is not possible, and instead, one needs to rely on alternative ways of solving the Schrödinger bridge. The two main methods to do so are given by iterative proportional fitting (IPF, De Bortoli et al.,, 2021) and Schrödinger bridge matching (SBM, Shi et al.,, 2023) which we review below.
@@ A.2.1 Conditional score matching
Score-based generative models can readily generate conditional samples from pre-trained unconditional models by plugging $\nabla_{x}\log\pi(y\mid x)$ into the drift of (9). However, it is in general hard to know $\pi(y\mid x)$ or its score, for example, in image restoration/classification applications (Song et al.,, 2021). Due to this, one often needs to approximate a separate time-varying model for $\pi_{t}(y\mid x_{t})$ or make domain-based heuristics, but this inevitably incurs additional errors for the conditional sampling.
@@ A.2.2 Conditional Schrödinger bridges
Song et al., (2021) which requires learning or approximating the likelihood $\pi_{t}(y\mid x_{t})$. By contrast, Shi et al., (2022) propose a method to learn a conditional sampler only requiring being able to sample artificial observations $\pi(y\mid x)$. The idea is built upon constructing a Schrödinger bridge (see Section A.1.2) for the augmented distribution $\pi(x,y)$, namely
@@
with initial distribution $X_{0},Y_{0}\sim\pi(x,y)$ and an arbitrary drift $a$. In Equation (14), $\pi_{\mathrm{ref}}(x\mid y)$ refers to any informative approximation to $\pi(x\mid y)$ that is easy to sample from. We can then apply the methodology introduced in Section A.1.2 to solve Equation (14).
@@ A.3 Twisted diffusion samplers
for which the conditional distribution of $U_{T}\mid y$ is $\pi(\cdot\mid y)$. This corresponds to a state-space model with a final observation $Y=y$, which can therefore be sampled by running the reverse diffusion (16) and then correcting the samples via importance sampling at the final time step. This is however well-known to be inefficient (Whiteley and Lee,, 2014) as the unconditional samples $U_{T}$ are unlikely to be distributed according to $\pi(\cdot\mid y_{T})$.
In order to efficiently sample from $\pi(\cdot\mid y)$, Wu et al., (2023) proposed to form the twisted model
@@
which marginally recovers $\pi(U_{T}\mid y)$. Once an approximation $\tilde{p}(u_{t_{k+1}}\mid y,u_{t_{k}})$ to the true conditional process $p(u_{t_{k+1}}\mid y,u_{t_{k}})$ is chosen, the twisted model (17) can then be sampled from using particle filtering, with a proposal $\tilde{p}(u_{t_{k+1}}\mid y,u_{t_{k}})$ and the corresponding importance weights as given in (18), see also Wu et al., (2023); Whiteley and Lee, (2014) for more details. While principled, this approach presents at least the following two issues: (i) it requires the likelihood model $\pi(y\mid x)$ to be known, which is not the case in most applications of interest, (ii) and it is computationally intensive as forming the approximation $\tilde{p}$ requires the denoising score function to be differentiated (Wu et al.,, 2023, Section 3.2).
@@ A.4 Pathwise conditioning
The idea of conditioning on the path of a diffusion process has been explored in the context of diffusion models by Trippe et al., (2023) for scaffolding and by Dou and Song, (2024, concurrent with our paper) for the slighly more general setting of inverse problems. The idea is that, given an observation $Y=y$, and provided that the forward noising diffusion for $X_{t},Y_{t}$ is separable, one can form a sequence of observations $Y_{0:K}$ and then sample from the conditional distribution $\pi(X_{0:K}\mid Y_{0:K})$ by running the reverse diffusion (5) conditioned on $Y_{0:K}$ for instance by using a particle filter as in Wu et al., (2023). Beside the need for a separable noising SDE, the method is asymptotically consistent under three assumptions:
@@
Appendix B Particle Markov chain Monte Carlo
Particle Markov chain Monte Carlo (Andrieu et al.,, 2010, PMCMC) is a class of Markov chain Monte Carlo (MCMC) algorithms targetting the posterior distribution of a state-space model $\pi(x_{0:T})\propto p(x_{0})\prod_{t=1}^{T}p(x_{t}\mid x_{t-1})\prod_{t=0}^{T}p(y_{t}\mid x_{t})$, where $y_{t}$ is a given set of observations.
Two different versions of PMCMC exist: the first one, known as the particle independent Metropolis–Hastings (PIMH) algorithm, is a pseudo-marginal MCMC (Andrieu and Roberts,, 2009) algorithm that uses a particle filter to both propose a full new trajectory, and estimate the likelihood of the observations given the state, and the second one, known as the conditional SMC (cSMC) algorithm, is a Gibbs-like ensemble sampler (Tjelmeland,, 2004) updates a trajectory sequentially using a particle filter.
Both algorithms can also be used to perform parameter inference in state-space models, by targeting the joint posterior distribution of the parameters and the states. The corresponding instances of the algorithms are known as the particle marginal Metropolis–Hastings (PMMH) algorithm and the particle Gibbs (pGibbs) algorithm, respectively. This is particularly useful in our context, where the observation trajectory $y_{1:T}$, where $y_{0}$ is our data, plays the role of a parameter in the state-space model with prior $p(y_{1:T}\mid y_{0})$ being given by the forward noising diffusion model (5).
Give the current state of the Markov chain $X^{n}_{0:T},Y^{n}_{1:T}$, the PMMH algorithm proceeds as follows:
  1. 1.
    Propose a new trajectory $y_{1:T}\sim q(y_{1:T}\mid y^{n}_{1:T})$;
  2. 2.
    Sample a new trajectory $x_{0:T}\sim\pi(x_{0:T}\mid y_{1:T})$ and the corresponding normalising constant estimator $\hat{\pi}(y_{1:T})$ from a particle filter;
  3. 3.
    Compute the acceptance probability of the new state and observations $(x_{0:T},y_{1:T})$ as
    $\alpha=\frac{\hat{\pi}(y_{1:T})q(y^{n}_{1:T}\mid y_{1:T})p(y_{1:T}\mid y_{0})}{\hat{\pi}(y^{n}_{1:T})q(y_{1:T}\mid y^{n}_{1:T})p(y^{n}_{1:T}\mid y_{0})},$ (19)
  4. 4.
    Accept the new pair $(x_{0:T},y_{1:T})$ with probability $\min\left\{1,\alpha\right\}$.
In step 2, the trajectory is obtained as a fixed-lag smoothing estimate of the state trajectory (Chopin and Papaspiliopoulos,, 2020, Chap. 12) and in step 19, the normalising constant estimator $\hat{\pi}(y^{n}_{1:T})$ is the one that was computed when the current state $(x^{n}_{0:T},y^{n}_{1:T})$ was accepted.
On the other hand, the pGibbs algorithm proceeds as follows:
  1. 1.
    Sample a new trajectory $y^{n+1}_{1:T}\sim\pi(y_{1:T}\mid x^{n}_{0:T},y_{0})$;
  2. 2.
    Sample a new trajectory $x^{n}_{0:T}\sim K(\cdot\mid y_{1:T},x^{n}_{0:T})$ from the so-called conditional SMC kernel $K$, given in Algorithm 1.
While the full description of cSMC is given in Algorithm 1, we provide a simplified version of the method in Algorithm 2 that considers a single time step, thereby recovering algorithms from Tjelmeland, (2004).
Algorithm 2 Conditional Sequential Monte Carlo (cSMC) for a single time step
0:  $x_{0}$, $y_{0}$
0:  A new $x_{0}$
  for $n=2,\ldots,N$ do
     Sample $x_{0}^{n}\sim p(x_{0})$
  Set $x_{0}^{1}=x_{0}$
  for $n=1,\ldots,N$ do
     Compute $w^{n}=p(y_{0}\mid x_{0}^{n})$
  Sample $m\in\{1,\ldots,N\}$ with probabilities proportional to $w^{1:N}$
  Set $x_{0}=x_{0}^{m}$
In practice, PMMH and PGibbs have different benefits and drawbacks: PMMH allows to sample the parameter $y_{1:T}$ directly, but is hindered by the variance of the likelihood estimator, while PGibbs is more stable but requires to sample the parameter $y_{1:T}$ conditionally on the state $x_{0:T}$, which can be inefficient when the two are highly correlated. Notwithstanding, CSMC is usually preferred in practice because, contrary to PMMH, it recovers a perfect Gibbs sampler when $N\to\infty$, while the acceptance rate of PMMH will usually not converge to one, even in the limit of infinite particles (Andrieu et al.,, 2018, Appendix).
Appendix C Conditional sampling under linear observation models
Concurrently to our work, Dou and Song, (2024) have extended Trippe et al., (2023)’s work to the case of linear observation models, whereby they are given a diffusion model targeting the unconditional distribution $\pi(x)$, and a linear observation model $y=Ax+c\xi$, where $A$ is a linear operator, $\xi$ is a standard Gaussian noise, and $c>0$ is the observation noise scale. Then they aim to sample from the conditional distribution $\pi(x\mid y)$.
Assuming that the forward noising SDE and its reverse denoising
$\begin{split}\mathop{}\!\mathrm{d}X_{t}=-\frac{\sigma^{2}}{2}X_{t}\mathop{}\!\mathrm{d}t+\sigma\mathop{}\!\mathrm{d}W^{X}_{t},\quad\mathop{}\!\mathrm{d}U_{t}=\mu^{U}(U_{t},t)\mathop{}\!\mathrm{d}t+\sigma\mathop{}\!\mathrm{d}W^{U}_{t},\end{split}$ (20)
are known [^3] Here we have assumed that the noising SDE had constant diffusion for the sake of ease of exposition., they proceed to sample from $\pi(x\mid y)$ by forming an augmented noising diffusion model for the joint distribution of $x_{[0,T]}$ and $y_{(0,T]}$ conditioned on $y_{0}$:
$\begin{split}\mathop{}\!\mathrm{d}Y_{t}=-\frac{\sigma^{2}}{2}Y_{t}\mathop{}\!\mathrm{d}t+\sigma A\mathop{}\!\mathrm{d}W_{t},\qquad\mathop{}\!\mathrm{d}X_{t}=-\frac{\sigma^{2}}{2}X_{t}\mathop{}\!\mathrm{d}t+\sigma\mathop{}\!\mathrm{d}W_{t},\end{split}$ (21)
for the same noise process $W_{t}$ for both $X$ and $Y$ and $(X_{0},Y_{0})\sim\pi(x)\mathcal{N}(y;Ax,c^{2}I)$
They then show that the reverse pair of SDEs follows, after discretisation:
$(u_{0:K},v_{0:K})\sim\mathbb{B}(u_{0:K})\prod_{k=0}^{K}\mathcal{N}(v_{k};Au_{k},c^{2}\lambda_{k}I),$ (22)
where the variance factor $\lambda_{k}$ can be computed in closed form. As a consequence, once $(Y_{k})_{k=0}^{K}$ has been obtained from (21), they can obtain approximate samples for $U_{0:K}\mid\{V_{0:K}=Y_{K:0}\}$ by using a particle filter. This approach is highly similar to Trippe et al., (2023): it is a biased pseudo-marginal algorithm, and suffers from the same bias as they do. Consequently, we can correct their approach in exactly the same way as we have corrected the method of Trippe et al., (2023) in this article: rather than implementing an approximate sampler for $\mathbb{B}(u_{0:K}\mid V_{0:K})$, we can implement a Markov kernel keeping it invariant such as the CSMC or correct pseudo-marginal kernels we propose in this article.
Appendix D Correction of the measurement path
One minor caveat of our Gibbs-CSMC sampler is that the path $Y_{1:K}$ was not drawn (marginally) from the backward model $\mathbb{B}$ but from $\mathbb{F}$. Specifically, since the learnt reversal is almost always an estimated one in practice, the path $V_{K:1}=Y_{1:K}$ drawn from the forward process $\mathbb{B}$ is not a sample of the reversal $\mathbb{B}$, resulting in a statistical bias. However, we show that (i) this bias can be fully corrected with an intermediary Metropolis–Hasting step, and that (ii) this bias empirically does not impact the results when ignored. First, we show how to remove the bias. Define $Y_{1:K}^{*}$ as the proposal path sampled from the forward process, and $Y_{1:K}^{j}$ as the previous Gibbs chain sample (the same goes for $X_{1:K}^{*}$ and $X_{1:K}^{j}$), we accept/reject the proposal based on the probability
$\min\Biggl{(}1,\frac{q(Y_{1:K}^{*}{\;|\;}X_{0:K}^{*},y_{0})\,p(Y_{1:K}^{j},X_{1:K}^{j}{\;|\;}X_{0}^{j},y_{0})}{q(Y_{1:K}^{j}{\;|\;}X_{0:K}^{j},y_{0})\,p(Y_{1:K}^{*},X_{1:K}^{*}{\;|\;}X_{0}^{*},y_{0})}\Biggr{)},$ (23)
where $q$ is the (approximate) backward process density, and $p$ is the (exact) forward density. This kernel then targets the correct distribution with the approximate prior diffusion. Lastly, we perform a new experiment to verify the bias of the mismatch (when ignored) by extending the MNIST super-resolution experiment. We remark that this experiment is on a Schrödinger bridge model, where the issue would be more pronounced since both the forward and backward are learnt, and, therefore are more likely to be inconsistent. We use the identical experiment setting as in Section 4.2, except that now we involve the MH kernel correction of Equation (23) in the Gibbs-CSMC sampler (using 10 particles). The results are shown in Table 4, and we conclude that for this experiment the bias is not significant, and in practice, the MH correction may indeed be unnecessary.
Table 4: The effect of correcting the measurement path in Section D. The numbers are averaged over 100 independent runs and 100 MC samples. We see that Gibbs-CSMC with correction has hardly differs from its uncorrected version, showing that the bias is empirically negligible.
Method PSNR (std) SSIM (std) MH acceptance ratio (std)
Gibbs-CSMC without correction 10.6373 (2.0228) 0.2806 (0.1368) 1.00 (0.00)
Gibbs-CSMC with MH correction 10.6370 (2.0099) 0.2820 (0.1382) 0.48 (0.50)
Appendix E Experiment settings